# Subspace Ordering for Maximum Response Preservation in Sufficient Dimension Reduction

### v1

@@ Abstract
Sufficient dimension reduction (SDR) methods aim to identify a dimension reduction subspace (DRS) that preserves all the information about the conditional distribution of a response given its predictor. Traditional SDR methods determine the DRS by solving a method-specific generalized eigenvalue problem and selecting the eigenvectors corresponding to the largest eigenvalues. In this article, we argue against the long-standing convention of using eigenvalues as the measure of subspace importance and propose alternative ordering criteria that directly assess the predictive relevance of each subspace. For a binary response, we introduce a subspace ordering criterion based on the absolute value of the independent Student’s T-statistic. Theoretically, our criterion identifies subspaces that achieve the local minimum Bayes’ error rate and yields consistent ordering of directions under mild regularity conditions. Additionally, we employ an F-statistic to provide a framework that unifies categorical and continuous responses under a single subspace criterion. We evaluate our proposed criteria within multiple SDR methods through extensive simulation studies and applications to real-data. Our empirical results demonstrate the efficacy of reordering subspaces using our proposed criteria, which generally improves classification accuracy and subspace estimation compared to ordering by eigenvalues.
@@ 1 Introduction
Most classical SDR methods rely on slicing the data into $H$ contiguous non-overlapping intervals to construct functions of the first two conditional moments with the goal of recovering the CS. Notable examples of this approach include sliced inverse regression (SIR), introduced by Li, (1991), and sliced average variance estimation (SAVE), proposed by Cook and Weisberg, (1991). This slicing-based framework is particularly suited to a continuous response, where slicing facilitates tractable estimation of conditional expectations and yields a DRS that is contained in $\mathcal{S}_{Y|\mathbf{X}}$ for sufficiently large $H$. When the response is categorical, however, slicing becomes trivial since the grouping of observations is explicit by the $H$ distinct populations. Moreover, with a categorical response, the emphasis is often on optimizing a classification rule rather than examining other aspects of $Y|\mathbf{X}$. Thus, Cook and Yin, (2001) proposed the central discriminant subspace (CDS) as an alternative target subspace to the CS for discriminant analysis. Let the Bayes’ rule be $\phi(\mathbf{X})\coloneqq\arg\max_{h=1,\ldots,H}\Pr\left(Y=h|\mathbf{X}\right)$. For a subspace $\mathcal{S}\subseteq\mathbb{R}^{p}$, let $\phi_{\mathcal{S}}(\mathbf{X})\coloneqq\arg\max_{h=1,\ldots,H}\Pr\left(Y=h|\mathbf{P}_{\mathcal{S}}\mathbf{X}\right)$. Moreover, for any basis matrix $\boldsymbol{\beta}$ such that $\text{span}(\boldsymbol{\beta})=\mathcal{S}$, we have $\phi_{\mathcal{S}}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\Pr\left(Y=h|\boldsymbol{\beta}^{\top}\mathbf{X}\right)$. Thus, if $\mathcal{S}$ satisfies $\phi_{\mathcal{S}}(\mathbf{X})=\phi(\mathbf{X})$, then it is a discriminant subspace. The CDS, denoted as $\mathcal{S}_{D(Y|\mathbf{X})}\subseteq\mathcal{S}_{Y|\mathbf{X}}$, is then defined as the intersection of all discriminant subspaces, given that the intersection itself is a discriminant subspace.
@@
More specifically, when the response is binary, we propose using the absolute value of the independent Student’s T-statistic, introduced by Welch, (1947), as a simple yet effective importance measure for a DRS. We naturally extend the idea of a subspace ordering criterion to a categorical response with more than two populations by employing an F-statistic, introduced by Fisher, (1925). Furthermore, within the slicing-based framework, we demonstrate that an F-statistic can also serve as a subspace criterion in the continuous response setting. That is, by emphasizing maximum slice separation, we can often identify subspaces that best capture the conditional moments. Although methods exist to determine the structural dimension $d$ of the CS (e.g., see Chapters 9 and 10 in Li, (2018) and see Zeng et al., (2024)), to the best of our knowledge, we are the first to propose reordering the DRS by a criterion different from the eigenvalue magnitudes. We believe that, while the intuition of these proposed criteria is simple, its simplicity offers a novel and interpretable importance measure for subspaces that addresses a previously overlooked but crucial aspect of subspace selection—namely, the need to directly assess predictive relevance rather than defaulting to eigenvalue magnitude.
@@ 2.2 Discriminant Analysis
$\phi_{QDA}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\left\{\log\pi_{h}+\frac{1}{2}\log|\boldsymbol{\Sigma}_{h}|-\frac{1}{2}(\mathbf{X}-\boldsymbol{\mu}_{h})^{\top}\boldsymbol{\Sigma}_{h}^{-1}(\mathbf{X}-\boldsymbol{\mu}_{h})\right\},$ (2)
@@
$\phi_{LDA}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\left\{(\boldsymbol{\mu}_{h}-\boldsymbol{\mu}_{1})^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{X}-[\boldsymbol{\mu}_{h}-\boldsymbol{\mu}_{1}]/2)\right\}.$ (6)
@@
where $\Phi(\cdot)$ is the cumulative distribution function (CDF) of a standard normal random variable. This expression yields the optimal Bayes’ error rate based on the one-dimensional (1D) projection of $\mathbf{X}\in\mathbb{R}^{p}$ onto the direction $\boldsymbol{\Sigma}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})$, which spans $\mathcal{B}$. Thus, by Lemma 2, we have $\mathcal{B}=\mathcal{S}_{D(Y\mid\mathbf{X})}$ under the LDA rule, which confirms that maximum population separation, and consequently the OER, is achieved when we project observations onto the CDS. Therefore, for any subspace spanned by an arbitrary nonzero vector $\mathbf{v}\in\mathbb{R}^{p}$, the Bayes’ error rate in that subspace will exceed the global OER in (8) unless $\mathbf{v}$ spans the same subspace as $\mathcal{B}$. However, among a set of $p$ candidate directions $\mathbf{v}_{1},\ldots,\mathbf{v}_{p}$, we can use the Bayes’ OER to determine the subspace most aligned with $\mathcal{S}_{D(Y\mid\mathbf{X})}$. Thus, under model (1), we have $\mathbf{v}^{\top}_{j}\mathbf{X}|\{Y=h\}\sim\mathcal{N}(\mathbf{v}_{j}^{\top}\boldsymbol{\mu}_{h},\mathbf{v}^{\top}_{j}\boldsymbol{\Sigma}_{h}\mathbf{v}_{j})$, and, consequently, the 1D Bayes’ OER in the subspace defined by $\text{span}(\mathbf{v}_{j})$ is given by
@@ 2.3 Select SDR Methods
SIR-II $\mathbb{E}\left[\text{Var}(\mathbf{Z}|Y)-\mathbb{E}\{\text{Var}(\mathbf{Z}|Y)\}\right]^{2}$ $\boldsymbol{\Sigma}_{\mathbf{X}}$
DR $\big\{\frac{1}{2}[2\mathbb{E}[\mathbb{E}^{2}(\mathbf{ZZ}^{T}\mid Y)]+2\mathbb{E}^{2}[\mathbb{E}(\mathbf{Z}\mid Y)\mathbb{E}(Z^{T}\mid Y)]+$ $\boldsymbol{\Sigma}_{\mathbf{X}}$
$2\mathbb{E}\left[\mathbb{E}(\mathbf{Z}\mid Y)\mathbb{E}(\mathbf{Z}\mid Y)\right]\mathbb{E}\left[\mathbb{E}(\mathbf{Z}\mid Y)\mathbb{E}(\mathbf{Z}^{T}\mid Y)\right]-2\mathbf{I}_{p}]\}^{1/2}$
@@
where $\mathbf{M}\in\mathbb{R}^{p\times p}$ is a method-specific symmetric kernel matrix, and $\mathbf{N}\in\mathbb{S}^{p}_{+}$ is often taken to be the common covariance matrix, denoted as $\boldsymbol{\Sigma}_{\mathbf{X}}$. Additionally, $\mathbf{v}_{1},\ldots,\mathbf{v}_{p}$ are the eigenvectors satisfying $\mathbf{v}_{j}^{\top}\mathbf{N}\mathbf{v}_{\ell}=1$ for $j=\ell$ and $\mathbf{v}_{j}^{\top}\mathbf{N}\mathbf{v}_{\ell}=0$ for $\ell\neq j$, corresponding to the eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{p}$. We denote the generalized eigenvalue problem with matrices $\mathbf{M}$ and $\mathbf{N}$ as $\text{GEV}(\mathbf{M},\mathbf{N})$. Let $\mathbf{Z}\coloneqq\boldsymbol{\Sigma}_{\mathbf{X}}^{-1/2}\{\mathbf{X}-\mathbb{E}(\mathbf{X})\}$. Then, we summarize the generalized eigenvalue problem for the following SDR methods in Table 1: principal components analysis (PCA) introduced by Pearson, (1901), SIR by Li, (1991), SAVE by Cook and Weisberg, (1991), a variant of SIR that considers second-order moments referred to as sliced inverse regression II (SIR-II) by Li, (1991), directional regression (DR) by Li and Wang, (2007), which aims to achieve an exhaustive estimate of the CS, and stabilized sufficient dimension reduction (SSDR) by Boonstra et al., (2025), which adjusts for heteroscedasticity and estimates a similar subspace as the CDS under QDA.
@@ 3 Subspace Ordering Criterion for Binary Response Preservation
3.1 Illustrative Example of Potential Gain by Reordering Subspaces
@@
For each direction $\mathbf{v}_{j}$, we define the projected population SNR between populations as
@@
To illustrate $T_{j}$ as a subspace criterion, we consider a simple example in which the leading eigenvectors do not necessarily correspond to the subspace that best preserves the relationship between the response and predictors. For this example, we focus on the simple SDR method of PCA. From Table 1, PCA can be formulated as solving $\text{GEV}(\boldsymbol{\Sigma}_{X},\mathbf{I}_{p})$, and the resulting DRS is defined by $\text{span}(\boldsymbol{\Sigma}_{X})$. Consider the covariance structure $\boldsymbol{\Sigma}_{X}=\text{diag}(3,2,1)$, a diagonal matrix with diagonal entries 3, 2, and 1. One can easily show that the eigenvectors of $\boldsymbol{\Sigma}_{X}$ are $\mathbf{v}_{1}=(1,0,0)^{\top}$, $\mathbf{v}_{2}=(0,1,0)^{\top}$, and $\mathbf{v}_{3}=(0,0,1)^{\top}$, and the respective eigenvalues are $\lambda_{1}=3$, $\lambda_{2}=2$, and $\lambda_{3}=1$. Thus, because $\lambda_{1}>\lambda_{2}>\lambda_{3}$, the basis for the traditional DRS is given by $\mathbf{V}=(\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3})$, with $\mathbf{v}_{1}$ corresponding to the first DRS with the largest variance. However, as we will demonstrate, $\mathbf{v}_{1}$ does not necessarily span the subspace that best captures the relationship between the response and the predictors. In particular, when we consider a binary response, we show that $T_{j}$ provides a better criterion for maximum population separation.
@@ 4 Subspace Criterion for a Categorical or Continuous Response
When the response $Y$ is discretized, either naturally through predefined populations or artificially through slicing, such that $\tilde{Y}\in\{1,\ldots,H\}$ with $H\geq 2$, we formalize $F_{j}$ as a subspace criterion by studying the theoretical properties of its population analogue,
@@
Let $\mathbf{X}\in\mathbb{R}^{p}$ and $\tilde{Y}\in\{1,\ldots,H\}$. Let $\boldsymbol{\Sigma}=\sum_{h=1}^{H}\pi_{h}\boldsymbol{\Sigma_{h}}$ and $\Psi_{j}$ be as defined in (15). We then have the following results.
@@
    Suppose $\mathbf{X}|\tilde{Y}$ follows model (1). If $\Psi_{j}\neq 0$, then $\boldsymbol{\Sigma}\mathbf{v}_{j}\not\perp\mathcal{S}_{D(Y\mid\mathbf{X})}$.
@@
Theorems 4 and 5 establish $F_{j}$ as a consistent sample-based subspace criterion that preserves the ranking induced by $\Psi_{j}$, thereby providing a measure for evaluating the relevance of candidate directions. We note that, in contrast to the binary case where $\Delta_{j}$ yields a direct monotonic relationship with the Bayes error rate, such a result is not available for $\Psi_{j}$ because no general tractable OER for LDA or QDA exists when $H>2$. However, our simulation results and real-data applications demonstrate the practical effectiveness of ordering subspaces by $F_{j}$, which identifies directions with greater overall class separation and, as a result, often significantly reduces the estimated Bayes’ error rate.

### v2

@@ Abstract
Sufficient dimension reduction (SDR) methods aim to identify a dimension reduction subspace (DRS) that preserves all the information about the conditional distribution of a response given its predictor. Traditional SDR methods determine the DRS by solving a method-specific generalized eigenvalue problem and selecting the eigenvectors corresponding to the largest eigenvalues. In this article, we argue against the long-standing convention of using eigenvalues as the measure of subspace importance and propose alternative ordering criteria that directly assess the predictive relevance of each subspace. For a binary response, we introduce a subspace ordering criterion based on the absolute value of the independent Student’s T-statistic. Theoretically, our criterion identifies subspaces that achieve the local minimum Bayes’ error rate and yields consistent ordering of directions under mild regularity conditions. Additionally, we employ an F-statistic to provide a framework that unifies categorical and continuous responses under a single subspace criterion. We evaluate our proposed criteria within multiple SDR methods through extensive simulation studies and applications to real-data. Our empirical results demonstrate the efficacy of reordering subspaces using our proposed criteria, which generally yields significant improvements in classification accuracy and subspace estimation compared to ordering by eigenvalues.
@@ 1 Introduction
Most classical SDR methods rely on slicing the data into $H$ contiguous non-overlapping intervals to construct functions of the first two conditional moments with the goal of recovering the CS. Notable examples of this approach include sliced inverse regression (SIR), introduced by Li, (1991), and sliced average variance estimation (SAVE), proposed by Cook and Weisberg, (1991). This slicing-based framework is particularly suited to a continuous response, where slicing facilitates tractable estimation of conditional expectations and yields a DRS that is contained in $\mathcal{S}_{Y|\mathbf{X}}$ for sufficiently large $H$. When the response is categorical, slicing becomes trivial since the grouping of observations is explicit by the $H$ distinct populations. Moreover, with a categorical response, the emphasis is often on optimizing a classification rule rather than examining other aspects of $Y|\mathbf{X}$. Thus, Cook and Yin, (2001) proposed the central discriminant subspace (CDS) as an alternative target subspace to the CS for discriminant analysis. Let the Bayes’ rule be $\phi(\mathbf{X})\coloneqq\arg\max_{h=1,\ldots,H}\Pr\left(Y=h|\mathbf{X}\right)$. For a subspace $\mathcal{S}\subseteq\mathbb{R}^{p}$, let $\phi_{\mathcal{S}}(\mathbf{X})\coloneqq\arg\max_{h=1,\ldots,H}\Pr\left(Y=h|\mathbf{P}_{\mathcal{S}}\mathbf{X}\right)$. Moreover, for any basis matrix $\boldsymbol{\beta}$ such that $\text{span}(\boldsymbol{\beta})=\mathcal{S}$, we have $\phi_{\mathcal{S}}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\Pr\left(Y=h|\boldsymbol{\beta}^{\top}\mathbf{X}\right)$. Thus, if $\mathcal{S}$ satisfies $\phi_{\mathcal{S}}(\mathbf{X})=\phi(\mathbf{X})$, then it is a discriminant subspace. The CDS, denoted as $\mathcal{S}_{D(Y|\mathbf{X})}\subseteq\mathcal{S}_{Y|\mathbf{X}}$, is then defined as the intersection of all discriminant subspaces, given that the intersection itself is a discriminant subspace.
@@
More specifically, when the response is binary, we propose using the absolute value of an independent Student’s T-statistic, introduced by Welch, (1947), as a simple yet effective importance measure for a DRS. We naturally extend the idea of a subspace ordering criterion to a categorical response with more than two populations by employing an F-statistic, introduced by Fisher, (1925). Furthermore, within the slicing-based framework, we demonstrate that an F-statistic can also serve as a subspace criterion in the continuous response setting. That is, by emphasizing maximum slice separation, we can often identify subspaces that best capture the conditional moments. Although methods exist to determine the structural dimension $d$ of the CS (e.g., see Chapters 9 and 10 in Li, (2018) and see Zeng et al., (2024)), to the best of our knowledge, we are the first to propose reordering the DRS by a criterion different from the eigenvalue magnitudes. We believe that, while the intuition of these proposed criteria is simple, its simplicity offers a novel and interpretable importance measure for subspaces that addresses a previously overlooked but crucial aspect of subspace selection—namely, the need to directly assess predictive relevance rather than defaulting to eigenvalue magnitude.
@@ 2.2 Discriminant Analysis
$\phi_{QDA}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\left\{\log\pi_{h}+\frac{1}{2}\log|\boldsymbol{\Sigma}_{h}|+\frac{1}{2}(\mathbf{X}-\boldsymbol{\mu}_{h})^{\top}\boldsymbol{\Sigma}_{h}^{-1}(\mathbf{X}-\boldsymbol{\mu}_{h})\right\},$ (2)
@@
$\phi_{LDA}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\left\{\log(\pi_{h}/\pi_{1})+(\boldsymbol{\mu}_{h}-\boldsymbol{\mu}_{1})^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{X}-[\boldsymbol{\mu}_{h}+\boldsymbol{\mu}_{1}]/2)\right\}.$ (6)
@@
where $\Phi(\cdot)$ is the cumulative distribution function (CDF) of a standard normal random variable. This expression yields the optimal Bayes’ error rate based on the one-dimensional (1D) projection of $\mathbf{X}\in\mathbb{R}^{p}$ onto the direction $\boldsymbol{\Sigma}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})$, which is in $\mathcal{B}$. Thus, by Lemma 2, we have $\mathcal{B}=\mathcal{S}_{D(Y\mid\mathbf{X})}$ under the LDA rule, which confirms that maximum population separation, and consequently the OER, is achieved when we project observations onto the CDS. Therefore, for any subspace spanned by an arbitrary nonzero vector $\mathbf{v}\in\mathbb{R}^{p}$, the Bayes’ error rate in that subspace will exceed the global OER in (8) unless $\mathbf{v}$ spans the same subspace as $\mathcal{B}$. However, among a set of $p$ candidate directions $\mathbf{v}_{1},\ldots,\mathbf{v}_{p}$, we can use the Bayes’ OER to determine the subspace most aligned with $\mathcal{S}_{D(Y\mid\mathbf{X})}$. Thus, under model (1), we have $\mathbf{v}^{\top}_{j}\mathbf{X}|\{Y=h\}\sim\mathcal{N}(\mathbf{v}_{j}^{\top}\boldsymbol{\mu}_{h},\mathbf{v}^{\top}_{j}\boldsymbol{\Sigma}_{h}\mathbf{v}_{j})$, and, consequently, the 1D Bayes’ OER in the subspace defined by $\text{span}(\mathbf{v}_{j})$ is given by
@@ 2.3 Select SDR Methods
SIR-II $\mathbb{E}\left[\text{Cov}(\mathbf{Z}|Y)-\mathbb{E}\{\text{Cov}(\mathbf{Z}|Y)\}\right]^{2}$ $\boldsymbol{\Sigma}_{\mathbf{X}}$
DR $\boldsymbol{\Sigma}^{1/2}_{\mathbf{X}}\big\{[2\mathbb{E}[\mathbb{E}^{2}(\mathbf{ZZ}^{T}\mid Y)]+2\mathbb{E}^{2}[\mathbb{E}(\mathbf{Z}\mid Y)\mathbb{E}(\mathbf{Z}^{T}\mid Y)]+$ $\boldsymbol{\Sigma}_{\mathbf{X}}$
$2\mathbb{E}\left[\mathbb{E}(\mathbf{Z}\mid Y)\mathbb{E}(\mathbf{Z}\mid Y)\right]\mathbb{E}\left[\mathbb{E}(\mathbf{Z}\mid Y)\mathbb{E}(\mathbf{Z}^{T}\mid Y)\right]-2\mathbf{I}_{p}]\big\}\boldsymbol{\Sigma}^{1/2}_{\mathbf{X}}$
@@
where $\mathbf{M}\in\mathbb{S}^{p\times p}$ is a method-specific symmetric kernel matrix, and $\mathbf{N}\in\mathbb{S}^{p}_{+}$ is often taken to be the common covariance matrix, denoted as $\boldsymbol{\Sigma}_{\mathbf{X}}$. Additionally, $\mathbf{v}_{1},\ldots,\mathbf{v}_{p}$ are the eigenvectors satisfying $\mathbf{v}_{j}^{\top}\mathbf{N}\mathbf{v}_{\ell}=1$ for $j=\ell$ and $\mathbf{v}_{j}^{\top}\mathbf{N}\mathbf{v}_{\ell}=0$ for $\ell\neq j$, corresponding to the eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{p}$. We denote the generalized eigenvalue problem with matrices $\mathbf{M}$ and $\mathbf{N}$ as $\text{GEV}(\mathbf{M},\mathbf{N})$. Let $\mathbf{Z}\coloneqq\boldsymbol{\Sigma}_{\mathbf{X}}^{-1/2}\{\mathbf{X}-\mathbb{E}(\mathbf{X})\}$. Then, we summarize the generalized eigenvalue problem for the following SDR methods in Table 1: principal components analysis (PCA) introduced by Pearson, (1901), SIR by Li, (1991), SAVE by Cook and Weisberg, (1991), a variant of SIR that considers second-order moments referred to as sliced inverse regression II (SIR-II) by Li, (1991), directional regression (DR) by Li and Wang, (2007), which aims to achieve an exhaustive estimate of the CS, and stabilized sufficient dimension reduction (SSDR) by Boonstra et al., (2025), which adjusts for heteroscedasticity and estimates a similar subspace as the CDS under QDA.
@@ 3 Subspace Ordering Criterion for Binary Response Preservation
3.1 Method
@@
For each direction $\mathbf{v}_{j}$, we define the projected SNR between populations as
@@
To illustrate $T_{j}$ as a subspace criterion, we consider a simple example in which the leading eigenvectors do not necessarily correspond to the subspace that best preserves the relationship between the response and predictors. For this example, we focus on the simple SDR method of PCA. From Table 1, PCA can be formulated as solving $\text{GEV}(\boldsymbol{\Sigma}_{X},\mathbf{I}_{p})$, and the resulting DRS is defined by $\text{span}(\boldsymbol{\Sigma}_{X})$. Consider the covariance structure $\boldsymbol{\Sigma}_{X}=\text{diag}(3,2,1)$, a diagonal matrix with diagonal entries 3, 2, and 1. One can easily show that the eigenvectors of $\boldsymbol{\Sigma}_{X}$ are $\mathbf{v}_{1}=(1,0,0)^{\top}$, $\mathbf{v}_{2}=(0,1,0)^{\top}$, and $\mathbf{v}_{3}=(0,0,1)^{\top}$, and the respective eigenvalues are $\lambda_{1}=3$, $\lambda_{2}=2$, and $\lambda_{3}=1$. Thus, because $\lambda_{1}>\lambda_{2}>\lambda_{3}$, the basis for the traditional DRS is given by $\mathbf{V}=(\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3})$, with $\mathbf{v}_{1}$ corresponding to the first DRS with the largest variance. However, as we will demonstrate, $\mathbf{v}_{1}$ does not necessarily span the subspace that best captures the relationship between the response and the predictors. In particular, when we consider a binary response, we show that $T_{j}$ provides a better criterion for maximum class separation.
@@ 4 Subspace Criterion for a Categorical or Continuous Response
When the response $Y$ is discretized, either naturally through predefined populations or artificially through slicing, such that $\widetilde{Y}\in\{1,\ldots,H\}$ with $H\geq 2$, we formalize $F_{j}$ as a subspace criterion by studying the theoretical properties of its population analogue,
@@
Let $\mathbf{X}\in\mathbb{R}^{p}$ and $\widetilde{Y}\in\{1,\ldots,H\}$. Let $\boldsymbol{\Sigma}=\sum_{h=1}^{H}\pi_{h}\boldsymbol{\Sigma_{h}}$ and $\Psi_{j}$ be as defined in (15). We then have the following results.
@@
    Suppose $\mathbf{X}|\widetilde{Y}$ follows model (1). If $\Psi_{j}\neq 0$, then $\boldsymbol{\Sigma}\mathbf{v}_{j}\not\perp\mathcal{S}_{D(Y\mid\mathbf{X})}$.
@@
Theorems 4 and 5 establish $F_{j}$ as a consistent sample-based subspace criterion that preserves the ranking induced by $\Psi_{j}$, thereby providing a measure for evaluating the relevance of candidate directions. We note that, in contrast to the binary case where $\Delta_{j}$ yields a direct monotonic relationship with the Bayes’ error rate, such a result is not available for $\Psi_{j}$ because no general tractable OER for LDA or QDA exists when $H>2$. However, our simulation results and real-data applications demonstrate the practical effectiveness of ordering subspaces by $F_{j}$, which identifies directions with greater overall class separation and, as a result, often significantly reduces the estimated misclassification rate.